# Do LLMs Adhere to Label Definitions? Examining Their Receptivity to External Label Definitions

### v2

@@ 1 Introduction
Label Definitions are considered as grounded statements that provide context on what an AI model needs to do upon receiving a query  (Mu et al., 2024; Deng et al., 2025). These definitions are considered as clues to disambiguate unclear labels, helping models perform their tasks more effectively  (Peskine et al., 2023; Kumar et al., 2023; Xie et al., 2023). However, whether models truly process and incorporate these definitions into their decision-making remains unclear.
@@ 2.1 Knowledge Conflict
Conflicts arise when a model’s pre-trained knowledge conflicts with newly provided definitions, potentially affecting classification performance and explanation quality (Xie et al., 2023). To assess the model’s reliance on external definitions versus its internalized knowledge, we introduce two scenarios.
@@
Performance and Explanation Evaluation: To assess how definition manipulation affects model performance, we use the Matthews Correlation Coefficient (MCC), a robust metric for binary classification that accounts for all four confusion matrix categories (Matthews, 1975). MCC ranges from -1 to +1, where +1 indicates perfect classification, 0 represents random performance, and -1 indicates complete disagreement between predictions and ground truth . Unlike accuracy or F1-score, MCC remains reliable even with imbalanced datasets.
@@ 4.2 Definitions Integration: Results and Analysis
The fact-checking task (wice) reveals a different pattern. While GPT-4 shows modest benefits from definitions (25% MCC improvement with adjusted definitions), all models demonstrate remarkably consistent F1 scores for explanation alignment regardless of definition condition—a striking contrast to the variable impacts seen in other metrics and tasks. This unusual stability aligns with findings from Lin et al. (Lin et al., 2024), suggesting that for evidence identification tasks, which require factual recall rather than conceptual understanding, models rely primarily on distributional patterns learned during pretraining rather than explicit definitional guidance.
@@ 5 Related Work
While recent advances in reasoning methods, including chain-of-thought (CoT) prompting Wang et al. (2024); Zhao et al. (2024); Wei et al. (2023), Progressive-Hint Prompting Zheng et al. (2024), and PromptWizard Agarwal et al. (2024), have improved model performance, they cannot distinguish between genuine expert knowledge incorporation and superficial alignment Greenblatt et al. (2024). Our work uniquely reveals how LLMs semantically process external definitions, offering insights into whether models truly incorporate or merely imitate expert knowledge.
Building on various prompt engineering techniques Wu et al. (2024); Zhang et al. (2022); Yang et al. (2024); Fernando et al. (2023) and explanation strategies Lampinen et al. (2022), our work provides the first systematic analysis of how LLMs interpret and integrate semantic definitions. This is particularly crucial for high-stakes applications like mental health and hate speech detection, where eliciting the model’s true behavior rather than superficial compliance is essential—an aspect previous methods haven’t fully explored.

### v3

@@ 1 Introduction
Label Definitions are considered as grounded statements that provide context on what an AI model needs to do upon receiving a query  (Mu et al., 2024; Deng et al., 2025). These definitions are considered as clues to disambiguate unclear labels, helping models perform their tasks more effectively  (Peskine et al., 2023; Kumar et al., 2023; Xie et al., ). However, whether models truly process and incorporate these definitions into their decision-making remains unclear.
@@ 2.1 Knowledge Conflict
Conflicts arise when a model’s pre-trained knowledge conflicts with newly provided definitions, potentially affecting classification performance and explanation quality (Xie et al., ). To assess the model’s reliance on external definitions versus its internalized knowledge, we introduce two scenarios.
@@
Performance and Explanation Evaluation: To assess how definition manipulation affects model performance, we use the Matthews Correlation Coefficient (MCC), a robust metric for binary classification that accounts for all four confusion matrix categories (Baldi et al., 2000). MCC ranges from -1 to +1, where +1 indicates perfect classification, 0 represents random performance, and -1 indicates complete disagreement between predictions and ground truth . Unlike accuracy or F1-score, MCC remains reliable even with imbalanced datasets.
@@ 4.2 Definitions Integration: Results and Analysis
The fact-checking task (wice) reveals a different pattern. While GPT-4 shows modest benefits from definitions (25% MCC improvement with adjusted definitions), all models demonstrate remarkably consistent F1 scores for explanation alignment regardless of definition condition, a striking contrast to the variable impacts seen in other metrics and tasks. The stable explanation F1 on wice across definition conditions is consistent with its evidence-selection format (supporting sentence indices), which relies more on claim–evidence matching than label semantics [^4] Minor wording/citation correction relative to the published proceedings version..
@@ 5 Related Work
While recent advances in reasoning methods, including chain-of-thought (CoT) prompting Wang et al. (2024); Zhao et al. (2024); Wei et al. (2023), Progressive-Hint Prompting Zheng et al. , and PromptWizard Agarwal et al. (2024), have improved model performance, they cannot distinguish between genuine expert knowledge incorporation and superficial alignment Greenblatt et al. (2024). Our work uniquely reveals how LLMs semantically process external definitions, offering insights into whether models truly incorporate or merely imitate expert knowledge.
Building on various prompt engineering techniques Wu et al. (2024); Zhang et al. ; Yang et al. ; Fernando et al. (2023) and explanation strategies Lampinen et al. (2022), our work provides the first systematic analysis of how LLMs interpret and integrate semantic definitions. This is particularly crucial for high-stakes applications like mental health and hate speech detection, where eliciting the model’s true behavior rather than superficial compliance is essential—an aspect previous methods haven’t fully explored.